# Next-Token Prediction Task Assumes Optimal Data Ordering for LLM Training in Proof Generation

### v1

@@ Abstract
In the field of large language model (LLM)-based proof generation, despite being trained on extensive corpora such as OpenWebMath and Arxiv, these models still exhibit only modest performance on proving tasks of moderate difficulty. We believe that this is partly due to the suboptimal order of each proof data used in training. Published proofs often follow a purely logical order, where each step logically proceeds from the previous steps based on the deductive rules. However, this order aims to facilitate the verification of the proof’s soundness, rather than to help people and models learn the discovery process of the proof. In proof generation, we argue that the optimal order for one training data sample occurs when the relevant intermediate supervision for a particular proof step in the proof is always positioned to the left of that proof step. We call such order the intuitively sequential order. We validate our claims using two tasks: intuitionistic propositional logic theorem-proving and digit multiplication. Our experiments verify the order effect and provide support for our explanations. We demonstrate that training is most effective when the proof is in the intuitively sequential order. Moreover, the order effect and the performance gap between models trained on different data orders are substantial – with an 11 percent improvement in proof success rate observed in the propositional logic theorem-proving task, between models trained on the optimal order compared to the worst order.
@@ 1 Introduction
Recent works have shown the potential of large language models (LLMs) to perform mathematical reasoning and proof generation in both natural language and formalized environments Yang et al. (2023a); Welleck et al. (2021); Lample et al. (2022); Mikuła et al. (2023); Wang et al. (2023).
Meanwhile, leading models consume far more data than a human could (datasets like OpenWebMath contain 14B tokens), and still perform suboptimally on reasoning benchmarks Paster et al. (2023); Azerbayev et al. (2023); Touvron et al. (2023); Dubey et al. (2024); Paster et al. (2023). This is also true for downstream reasoning tasks where LLMs need further fine-tuning Yang et al. (2023b); An et al. (2024). We identify this phenomenon as the training inefficiency problem.
One possible explanation for such low training efficiency in proof generation for LLMs is poor internal order within each training data sample in the math domain. For instance, consider the purely logical order that many textbook proofs follow, where each step logically proceeds from the previous steps based on the deductive rules. This order is widely adopted for presenting proofs publicly because it allows for straightforward verification of the proof’s correctness. However, from a proof discovery perspective, this order is not intuitive, i.e. not the order in which the proof is actually discovered. For human learners, the most intuitive order of proof steps is one in which all steps contributing to the derivation of a particular proof step are presented prior to it. We call such order the intuitively sequential order. These steps, which aid in the derivation of a given proof step, are referred to as intermediate supervision for that step.
@@
Consider the following concrete scenario where the intuitively sequential order is preferred over the purely logical order in human learning and model training. A researcher wants to prove a math statement. He identifies a set of lemmas useful for the proof. But in order to transform the statement into a form where these lemmas can be applied, several preliminary steps are required. A textbook proof usually presents the preliminary steps first and then applies the lemmas, following the purely logical order. However, for LLM training – and for human learning – the helper lemmas are easy to figure out after seeing the statement. They provide intuition and act as intermediate supervision for coming up with the preliminary steps. On the other hand, the theorem statement itself provides no hints for these preliminary steps. The lemmas should therefore precede these preliminary steps in the training data
@@
We give a simple illustrative example of the difference between the purely logical order and the intuitively sequential order in Figure 1. We also provide a graduate-level math example in Appendix A.1. In Figure 1, Step 3 serves as the intermediate supervision that assist to predict Step 2, which in turn serves as the intermediate supervision that assist to predict Step 1. Inexperienced students typically struggle to understand this proof in the purely logical order on their first read, as it is unclear why the problem statement should lead to Step 1. If they are instead taught to first apply the lemma as in Step 3, the learning process becomes faster. In practice, the confusion of suboptimal order can be overcome through repeated reflection. As for model training with next token prediction, without introducing more elaborate training objectives, the only way to achieve similar effect is to modify the order of the training data.
@@
Given that (i) proofs and reasoning are often presented in the order of logic rather than in the intuitively sequential order, and (ii) next-token prediction is the predominant method used for pre-training and fine-tuning cutting-edge autoregressive LLMs, we hypothesize that training inefficiency of LLMs in solving reasoning tasks is partly due to the suboptimal order of data during training. We claim that the order of each data sample affect model training efficiency significantly, which we referred to as the order effect, and the intuitively sequential order is the optimal order for fine-tuning models. We show that there are two causes for this order effect. Assume we have proof step $a$, and proof steps $\{a_{i}\}$ that serve as intermediate supervision toward $a$, and they occur in the order ``$a,\{a_{i}\}"$. The first cause is that when learning to predict $a$, the model will not directly learn to use the supervision $\{a_{i}\}$, because it does not look forward during training. The second cause is that this order also induces the model to fit the non-existent dependency of $\{a_{i}\}$ on $a$, which we call the spurious relation. Note that in reasoning tasks, intermediate supervision is usually another sequence of reasoning steps.
@@
To verify our claims, we conduct experiments on two downstream tasks: (1) 4-by-4 digit multiplication and, (2) intuitionistic propositional logic theorem-proving. We use the multiplication task to perform further analysis. Various metrics were used to measure model performance. The results demonstrate that models fine-tuned/trained-from-scratch with the intuitively sequential proofs outperformed those fine-tuned/trained-from-scratch with proofs in other orders.
@@
Our experiments also validate the two causes for the order effect we propose. Despite the theoretical understanding that LLMs learn joint distributions over data, where updates from later tokens might influence a model’s prediction on an earlier token – our experiments find no evidence that LLMs can look forward during learning (weird here put to the end). Our models do not learn to use future tokens to predict the current proof step. We also show experimental evidence in the multiplication task that models are negatively impacted by spurious step dependencies learned during training when it’s trained on data in suboptimal orders.
Our contributions are summarized as follows:
@@
  • We demonstrate that data order significantly affects LLM training efficiency on reasoning tasks, which we refer to as the order effect.
@@
  • We identify that the intuitively sequential order is the optimal order for next-token-prediction pre-training/finetuning, where the intermediate supervision is naturally placed on the left.
@@
  • Through the intuitionistic propositional logic theorem-proving task and 4-by-4 digit multiplication task, our experiment shows that this order effect is due to next-token prediction. Two causes are (1) LLMs do not look forward during training, (2) LLMs may learn spurious token dependencies if trained with data samples in suboptimal order.
@@ 2 Related Work
Problem of next-token prediction training
Existing research has exposed weaknesses in the next-token prediction mechanism. For example, Berglund et al. (2023) showed the Reversal Curse: a model trained on “A is B" does not know that “B is A". Golovneva et al. (2024) proposed to train a model both sequentially and reversely, using two tasks together to enhance performance. In contrast, our work demonstrates that optimal orders exist for next-token prediction learning, specifically by the rule of positioning intermediate supervision to the left. Bachmann and Nagarajan (2024) highlighted a similar toy problem known as the star-path problem, where initial errors substantially reduced the model’s ability to predict later tokens.
Order effect on LLM: Context
Several works have studied how disrupting order in prompts might affect model performance. Chen et al. (2024) showed that LLMs achieve the best performance on GSM8K Cobbe et al. (2021) when the premise order aligns with the context required in intermediate reasoning steps. Liu et al. (2024) discover the lost in-the-middle phenomenon in the long-context scenario: LLMs perform best when relevant information is placed at the start or end of the input context, and perform poorly when important information is located in the middle.
Order effect on LLM: Digits
For long-digit addition, prior works demonstrated that reversing numbers in training data can help LLMs achieve better performance on digit-addition tasks Lee et al. (2023); Zhou et al. (2023b, 2024). While their work only applies to digit computation since they employed digit-wise reversal, our work studies the positional role of intermediate supervision in general for the proof generation task.
@@ 3.1 Problem setting
Given algorithm $\mathbf{A}$ and input $r_{0}$ from the problem space, we assume:
@@
    Finiteness.  Algorithm $\mathbf{A}$ generates a finite sequence of steps $\mathbf{A}(r_{0})=r_{1},\mathbf{A}(\{r_{0},r_{1}\})=r_{2},...,\mathbf{A}(\{r_{i}\}_{i=0}^{n-1})=r_{n_{r_{0}}}$, where $r_{n_{r_{0}}}$ is the last step that $\mathbf{A}$ will generate when given $r_{0}$, which usually is the final answer.
@@
    Strongest right-to-left dependent order among all possible orders.  Each step $r_{i}$ highly depends on $r_{j}$ for $1<=j<i$, i.e. the probability for $\mathbf{A}(\{r_{i}\}_{i=0}^{j-1})=r_{j}$ is high for all $r_{0}$ and $j$. For every $r_{0}$, $\{r_{i}\}_{i=0}^{m}$ serves as intermediate supervision for models to predict $r_{m+1}$ for all $0<=m<n_{r_{0}}$ in the sense of Wies et al. (2022). We call this order the intuitively sequential order. All other orders of the sequence have a weaker right-to-left dependency than the aforementioned order.
Given a model $M_{\theta}$ with $\theta$ as weights of the model, our goal is to enable $M_{\theta}$ to imitate algorithm $\mathbf{A}$ through the next-token prediction task. Specifically, we aim to find weights $\theta_{0}$ such that $\theta_{0}=\operatorname*{arg\,min}_{\theta}\frac{1}{n_{r_{0}}}\Sigma_{m=1}^{n_{r_{0}}}\mathbb{E}_{r_{0}}|M_{\theta}(\{r_{i}\}_{i=1}^{m})-\mathbf{A}(\{r_{i}\}_{i=1}^{m})|$. To put it in another way, we want the model to predict the entire sequence $\{r_{i}\}_{i=1}^{n_{r_{0}}}$ given $r_{0}$ for all $r_{0}$. Importantly, the loss function does not exclusively focus on $r_{n_{r_{0}}}$ but also accounts for the intermediate steps $r_{m}$ for $m=1,...,r_{n_{r_{0}-1}}$ when computing loss. This scenario arises in practice in theorem-proving tasks, where we are interested in both the intermediate proofs and the final answer.
@@ 3.2 Causes of the order effect
  • Model cannot look forward: intermediate supervision fails to help if placed on the right.  The model will not be able to learn the dependency of $r_{n_{r_{0}}}$ on $\{r_{1},...,r_{n_{r_{0}}-1}\}$. Through the next-token prediction training loss, given $r_{i}\in r$, the model has no way to look forward and use $r_{j}$ with $j>i$ to help learn to predict $r_{i}$. The model would find it much harder to learn to predict $r_{n_{r_{0}}}$ directly from $r_{0}$, since this task is not split into sub-tasks like given $\{r_{0}\}$ predict $r_{1}$, given $\{r_{0},r_{1}\}$ predict $r_{2}$, and so on Wies et al. (2022).
@@ 4.1.1 Intuitionistic Propositional Logic Theorem-Proving
For the first dataset, we use theorem statements and proofs in intuitionistic propositional logic (PropL), as introduced in An et al. (2024). A proof consists of tactics and intermediate states, where a tactic refers to an actual proof step, and a state represents what remains to be proved after a tactic is applied. Both the statements and the proofs are written in Lean, a popular formalized environment for theorem-proving in the mathematical community de Moura et al. (2015). We have two types of PropL data. In the first type of data, all proofs follow the original sequential order (referred to as SEQ order). We note that SEQ order follows the intuitively sequential order: all the previous tactics and states help generate the next tactic and state. For the second type, the proofs are reversed (referred to as SER order), such that the last tactic and the last state become the first tactic and the first state, the second-to-last tactic and the second-to-last state become the second tactic and the second state, and so forth. Note that this dataset was published after the Llama-2 family was released so the model hasn’t seen this dataset during pre-training. See an example of a statement and a proof in Appendix A.2 for demonstration.
@@
We use an algorithm to generate the problems, the proofs, and the final answers of 4-digit by 4-digit multiplication. It has been shown that such problems, even prompted with “Let’s think step by step”, is challenging for cutting-edge models like GPT-4 Liu and Low (2023). We create three types of proofs with the same proof steps but in different orders, whose compositions are summarized in the table. The orders are listed from the optimal to the worst. Examples of all three types of data can be found in the Appendix A.2 for demonstration. Note that for the multiplication task, for proofs in the SEQ order, given a proof step, all previous steps serve as the natural intermediate supervision toward that step, hence making this dataset an ideal dataset to test our claim.
For training and testing, we first enumerated all 4-digit by 4-digit tuples. The list of tuples was then randomly shuffled. We used the first 1,000,000 tuples as our training number tuples. Together with the proofs, they formed our training dataset. We used the last 1,000 number tuples to construct our testing sets for model evaluation. During training process, a problem together with its proof and final answer form one training data sample.
@@
Table 1: Three Types of Data. See Appendix A.2 for examples.

### v2

@@ Abstract
In the field of large language model (LLM)-based proof generation, despite extensive training on large datasets such as ArXiv, LLMs still exhibit only modest performance on proving tasks of moderate difficulty. We believe that this is partly due to the widespread presence of suboptimal ordering within the data for each proof used in training. For example, published proofs often follow a purely logical order, where each step logically proceeds from the previous steps based on the deductive rules. This order is designed to facilitate the verification of the proof’s soundness, rather than to help people and models learn the discovery process of the proof. In proof generation, we argue that the optimal order for one training data sample occurs when the relevant intermediate supervision for a particular proof step in the proof is always positioned to the left of that proof step. We call such order the intuitively sequential order. We validate our claims using two tasks: intuitionistic propositional logic theorem-proving and digit multiplication. Our experiments verify the order effect and provide support for our explanations. We demonstrate that training is most effective when the proof is in the intuitively sequential order. Moreover, the order effect and the performance gap between models trained on different data orders can be substantial – with an 11 percent improvement in proof success rate observed in the propositional logic theorem-proving task, between models trained on the optimal order compared to the worst order. Lastly, we define a common type of order issue in advanced math proofs and find that 17.3 percent of theorems with nontrivial proofs in the first two chapters of a widely used graduate-level mathematics textbook suffer from this issue. A detailed list of those proofs is provided in the appendix.
@@ 1 Introduction
Recent works have shown the potential of LLMs to perform mathematical reasoning and proof generation in both natural language and formalized environments (Yang et al., 2023a; Welleck et al., 2021; Lample et al., 2022; Mikuła et al., 2023; Wang et al., 2023).
Meanwhile, leading models consume far more data than a human could (datasets like OpenWebMath contain 14B tokens), and still perform suboptimally on reasoning benchmarks (Paster et al., 2023; Azerbayev et al., 2023; Touvron et al., 2023; Dubey et al., 2024; Paster et al., 2023). This is also true for downstream reasoning tasks where LLMs need further fine-tuning (Yang et al., 2023b; An et al., 2024). We identify this phenomenon as the training inefficiency problem.
One possible explanation for such low training efficiency in proof generation for LLMs is poor internal order within each training data sample in the math domain. For instance, consider the purely logical order that many textbook proofs follow, where each step logically proceeds from the previous steps based on the deductive rules. This order is widely adopted for presenting proofs publicly because it allows for straightforward verification of the proof’s correctness. However, from a proof discovery perspective, this order is not intuitive, i.e. not the order in which the proof is actually discovered. For human learners, the most intuitive order of proof steps is one in which all steps contributing to the discovery of a particular proof step are presented prior to it. We call such order the intuitively sequential order. These steps, which aid in the discovery of a given proof step, are referred to as intermediate supervision for that step.
@@
Consider a concrete scenario where an intuitive sequential order is preferable to a purely logical one in learning and model training. When proving a theorem, a researcher quickly identifies a key lemma but must first perform preliminary steps to apply it. Textbooks typically present these steps first, following a purely logical order. However, in LLM training—and human learning—these steps are easier to derive after seeing the key lemma, which serves as intermediate supervision to these steps. Without the lemma, it is unclear why these preliminary steps are necessary. We refer to this issue as Delayed Insight Issue (DII)—a problem where crucial insights are introduced too late, disrupting intuition and understanding. To avoid DII, training data should present the lemma before the preliminary steps in such cases. DII represents one of the many causes of the difference between the purely logical order and the intuitively sequential order for advanced math proofs.
@@
We give a simple illustrative example of DII in Figure 1. In Figure 1, Step 3 serves as the intermediate supervision that assists in predicting Step 2, which in turn serves as the intermediate supervision that assists in predicting Step 1. Inexperienced students typically struggle to understand this proof in the purely logical order on their first read, as it is unclear why the problem statement should lead to Step 1. If they are instead taught to first apply the lemma as in Step 3, the learning process becomes faster. In practice, the confusion of suboptimal order can be overcome through repeated reflection for humans. As for model training with next token prediction, without introducing more elaborate training objectives, the only way to achieve a similar effect is to modify the order of the training data.
@@
Given that (i) proofs and reasoning are often presented in the suboptimal order, and (ii) next-token prediction is the predominant method used for pre-training and fine-tuning cutting-edge autoregressive LLMs, we hypothesize that the training inefficiency problem of LLMs in solving reasoning tasks is partly due to the suboptimal order of each proof’s data during training. We claim that the intuitively sequential order is the optimal order for fine-tuning models. We show that there are two causes for the order effect. Assume we have proof step $a$, and proof steps $\{a_{i}\}$ that serve as intermediate supervision toward $a$, and they occur in the order ``$a,\{a_{i}\}"$. The first cause is that when learning to predict $a$, the model will not directly learn to use the supervision $\{a_{i}\}$, because it does not look forward during training. The second cause is that this order also induces the model to fit the non-existent dependency of $\{a_{i}\}$ on $a$, which we call the spurious relation. Note that in reasoning tasks, intermediate supervision is usually another sequence of reasoning steps.
@@
To verify our claims, we conduct experiments on two downstream tasks: (1) 4-by-4 digit multiplication and, (2) intuitionistic propositional logic theorem-proving. Various metrics were employed to measure model performance. The results consistently show that models fine-tuned/trained-from-scratch on the intuitively sequential proofs outperformed those fine-tuned/trained-from-scratch on proofs in other orders.
@@
Our experiments also validate the two causes for the order effect we propose. We find that models do not learn to use future tokens to predict the current proof step. We also show evidence in the multiplication task that models are negatively impacted by spurious step dependencies learned during training when it’s trained on data in suboptimal orders.
To demonstrate the prevalence of ordering issues in mathematical texts, we analyze the first two chapters of a widely used graduate-level textbook in analysis, which contain a total of 57 theorems with nontrivial proofs. We find that 17.3 percent of these proofs exhibit DII. A detailed list of the affected proofs is provided in Appendix A.2.
Our main contributions are summarized as follows:
@@
  • We find that data order significantly affects LLM training efficiency on reasoning tasks, which we refer to as the order effect.
@@
  • We claim optimal order to be the intuitively sequential order, where intermediate supervision is placed on the left.
@@
  • Through the propositional logic theorem-proving task and the digit multiplication task, our experiments show that this order effect is due to next-token prediction. Two causes are (1) LLMs do not look forward during training, (2) LLMs may learn spurious token dependencies if trained with data samples in suboptimal order.
  • We show that the Delayed Insight Issue—a specific ordering problem—frequently arises in advanced mathematical proofs, as illustrated by 9 examples from a popular textbook listed in the appendix.
@@ 2 Related Work
Problem of next-token prediction training.  Existing research has exposed weaknesses in the next-token prediction mechanism. For example, Berglund et al. (2023) showed the Reversal Curse: a model trained on “A is B" does not know that “B is A". Golovneva et al. (2024) proposed to train a model both sequentially and reversely, using two tasks together to enhance performance. In contrast, our work demonstrates that optimal orders exist for next-token prediction learning, specifically by the rule of positioning intermediate supervision to the left. Bachmann and Nagarajan (2024) highlighted a similar toy problem known as the star-path problem, where initial errors substantially reduced the model’s ability to predict later tokens.
Order effect on LLM: context.  Several works have studied how disrupting order in prompts might affect model performance. Chen et al. (2024) showed that LLMs achieve the best performance on GSM8K when the premise order aligns with the context required in intermediate reasoning steps. Liu et al. (2024) discover the lost in-the-middle phenomenon in the long-context scenario: LLM performance is best when relevant information is placed at the beginning or the end of the input context, while the performance is worst when the LLM needs to utilize information in the middle.
Order effect on LLM: digits.  For long-digit addition, prior works demonstrated that reversing numbers in training data can help LLMs achieve better performance on digit-addition tasks (Lee et al., 2023; Zhou et al., 2023b, 2024). While their work only applies to digit computation since they employed digit-wise reversal, our work studies the positional role of intermediate supervision in general for the proof generation task.
@@ 3.1 Problem setting
Given a nondeterministic algorithm $\mathbf{A}$ and input $r_{0}$ from the problem space, we assume:
@@
    Finiteness.  Algorithm $\mathbf{A}$ generates a finite sequence of steps $\mathbf{A}(r_{0})=r_{1},\mathbf{A}(\{r_{0},r_{1}\})=r_{2},...,\mathbf{A}(\{r_{i}\}_{i=0}^{n-1})=r_{n_{r_{0}}}$, where $r_{n_{r_{0}}}$ is the last step that $\mathbf{A}$ will generate when given $r_{0}$, which usually is the final answer. Here $n_{r_{0}}$ is the position of the last token (final answer) generated by algorithm $\mathbf{A}$ with input $r_{0}$.
@@
    Intuitively sequential order.  Each step $r_{i}$ highly depends on $r_{j}$ for $1<=j<i$, i.e. the probability for $\mathbf{A}(\{r_{i}\}_{i=0}^{j-1})=r_{j}$ is high for all $j$. For every $r_{0}$, $\{r_{i}\}_{i=0}^{m}$ serves as intermediate supervision for models to predict $r_{m+1}$ for all $0<=m<n_{r_{0}}$. We call this order the intuitively sequential order. All other orders of the sequence have a weaker right-to-left dependency than the aforementioned order.
Given a model $M_{\theta}$ with $\theta$ as weights of the model, our goal is to enable $M_{\theta}$ to imitate algorithm $\mathbf{A}$ through the next-token prediction task. Specifically, we aim to find weights $\theta_{0}$ such that $\theta_{0}=\operatorname*{arg\,min}_{\theta}\frac{1}{n_{r_{0}}}\Sigma_{m=1}^{n_{r_{0}}}\mathbb{E}_{r_{0}}|M_{\theta}(\{r_{i}\}_{i=1}^{m})-\mathbf{A}(\{r_{i}\}_{i=1}^{m})|$. To put it in another way, we want the model to predict the entire sequence $\{r_{i}\}_{i=1}^{n_{r_{0}}}$ given $r_{0}$. Importantly, the loss function does not exclusively focus on $r_{n_{r_{0}}}$ but also accounts for the intermediate steps $r_{m}$ for $m=1,...,r_{n_{r_{0}-1}}$ when computing loss. This scenario arises in practice in theorem-proving tasks, where we are interested in both the intermediate proofs and the final answer.
@@ 3.2 Causes of the order effect
  • Model cannot look forward: intermediate supervision fails to help if placed on the right.  The model will not be able to learn the dependency of $r_{n_{r_{0}}}$ on $\{r_{1},...,r_{n_{r_{0}}-1}\}$. Through the next-token prediction training loss, given $r_{i}\in\mathbf{r}=\{r_{i}\}_{i=0}^{n_{r_{0}}}$, the model has no way to look forward and use $r_{j}$ with $j>i$ to help learn to predict $r_{i}$. The model would find it much harder to learn to predict $r_{n_{r_{0}}}$ directly from $r_{0}$, since this task is not split into sub-tasks like given $\{r_{0}\}$ predict $r_{1}$, given $\{r_{0},r_{1}\}$ predict $r_{2}$, and so on (Wies et al., 2022).
@@ 4.1.1 Intuitionistic propositional logic theorem-proving
For the first dataset, we use theorem statements and proofs in intuitionistic propositional logic (PropL), as introduced in An et al. (2024). A proof consists of tactics and intermediate states, where a tactic refers to an actual proof step, and a state represents what remains to be proved after a tactic is applied. Both the statements and the proofs are written in Lean, a popular formalized environment for theorem-proving in the mathematical community (de Moura et al., 2015). We have two types of PropL data. In the first type of data, all proofs follow the original sequential order (referred to as SEQ order, see Table 1). We note that SEQ order follows the intuitively sequential order: all the previous tactics and states help generate the next tactic and state. For the second type, the proofs are reversed (referred to as SER order), such that the last tactic and the last state become the first tactic and the first state, the second-to-last tactic and the second-to-last state become the second tactic and the second state, and so forth. Note that this dataset was published after the Llama-2 family was released so the model hasn’t seen this dataset during pre-training. See an example of a statement and a proof in Appendix A.4.
@@
We use an algorithm to generate the problems, the proofs, and the final answers of 4-digit by 4-digit multiplication. It has been shown that such problems, even prompted with “Let’s think step by step”, are challenging for cutting-edge models like GPT-4 (Liu and Low, 2023). We create three types of proofs with the same proof steps but in different orders, whose compositions are summarized in the table. Examples of all three types of data can be found in the Appendix A.4. Note that for the multiplication task, for proofs in the SEQ order, given a proof step, all previous steps serve as the natural intermediate supervision toward that step, hence making this dataset ideal for testing our claims.
For training and testing, we first enumerated all 4-digit by 4-digit tuples. The list of tuples was then randomly shuffled. We used the first 1,000,000 tuples as our training number tuples. Together with the proofs, they formed our training dataset. We used the last 1,000 tuples to construct our testing number tuples.
Table 1: Three Types of Data. See Appendix A.4 for examples.